# Measurement-aligned Flow for Inverse Problem

### v1

@@ Abstract
Diffusion models provide a powerful way to incorporate complex prior information for solving inverse problems. However, existing methods struggle to correctly incorporate guidance from conflicting signals in the prior and measurement, especially in the challenging setting of non-Gaussian or unknown noise. To bridge these gaps, we propose Measurement-Aligned Sampling (MAS), a novel framework for linear inverse problem solving that can more flexibly balance prior and measurement information. MAS unifies and extends existing approaches like DDNM and DAPS, and offers a new optimization perspective. MAS can generalize to handle known Gaussian noise, unknown or non-Gaussian noise types. Extensive experiments show that MAS consistently outperforms state-of-the-art methods across a range of tasks.
@@ 1 Introduction
Here ${\epsilon\in\mathbb{R}^{m}}$ represents measurement noise, $x_{0}\in\mathbb{R}^{d}$ is drawn from data distribution $\pi_{0}(x_{0})$, $\mathcal{H}:\mathbb{R}^{d}\mapsto\mathbb{R}^{m}$ is the measurement function, and $y\in\mathbb{R}^{m}$ denotes the degraded measurement or observed image. A useful motivating example is a high-resolution image $x_{0}$, with a noisy degraded image $y$ and known corruption process.
Pretrained diffusion or flow models provide a prior distribution $\pi_{0}(x_{0})$ that significantly facilitates the solution of inverse problems. DPS (chung2022diffusion, ) and $\Pi$GDM (song2023pseudoinverse, ) estimate conditional scores directly from the measurement model by leveraging score decomposition to guide each diffusion sampling iteration. In contrast, FPS (dou2024diffusion, ), DAPS (zhang2024improving, ), and MPGD (he2023manifold, ) align denoiser outputs directly with measurements, thereby avoiding backpropagation. However, these methods are not explicitly designed to handle high levels of Gaussian noise, as addressed by approaches like DDNM (wang2022zero, ) and DDRM (kawar2022denoising, ).
Moreover, the above approaches lack the ability to effectively regulate the ‘strength’ and ‘type’ of the alignment between measurements and model predictions. They also commonly assume a known forward operator $\mathcal{H}$ and Gaussian noise with known variance, i.e., $\epsilon\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I}_{m})$. In practical settings, noise frequently deviates from Gaussian assumptions—exhibiting characteristics like salt-and-pepper, periodic, or Poisson distributions—or is completely unknown. Additionally, the forward measurement operator may be uncertain or inaccurately specified. Effectively addressing inverse problems under these more general and realistic conditions remains an open and challenging research area.
@@
  • We propose Measurement-Aligned Sampling (MAS), a novel framework for solving linear inverse problems. MAS provides additional optimization perspectives and generalizes approaches such as DDNM and DAPS for linear inverse problems. Furthermore, our proposed “overshooting” technique achieves superior restoration quality compared to DDNM across various inverse problem scenarios.
  • We develop new techniques to handle Gaussian noise and unknown noise. Specifically, we extend strategies akin to DDNM$+$ to address noisy inverse problems with known Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{y}\mathbb{I}_{m})$. Moreover, our novel parameterization scheme allows us to effectively handle noisy inverse problems with unknown or non-Gaussian noise structures and even non-differentiable measurements, such as JPEG restoration, without requiring explicit knowledge of the forward operator or noise level. The comparison of method applicability across different inverse problems is shown in Table 1.
  • Our experimental results demonstrate state-of-the-art performance across a diverse set of tasks, see examples in Fig. 1 and experiments in Sec. 5.
@@
Inverse Problem Noise strength DDNM DDRM $\Pi$GDM DAPS RED-Diff MAS (ours)
@@ 2 Background
Notation. Our time notation follows diffusion literature where $t=T$ corresponds to the source distribution and $t=0$ corresponds to the target distribution, consistent with the use of pretrained diffusion models in our experiments. In contrast, flow-based models typically adopt the reverse, with $t=0$ as the source distribution.
2.1 Conditional Diffusion and flow models
@@
Given training dataset $\mathcal{D}=\{x_{0}^{i}\}_{i=1}^{N}$ from target distribution $\pi_{0}(x_{0})$, $x_{0}^{i}\in\mathbb{R}^{d}$, the goal of generative modeling is to draw new samples from $\pi_{0}$. In the context of conditional generation, suppose that we have data samples from a joint disribution $(x_{0}^{i},y)\sim\pi(x_{0},y)$, where $x_{0}$, $y$ are dependent, and $y$ could be class labels or text information, for example.

### v2

# Measurement-Aligned Sampling for Inverse Problems

@@ Abstract
Diffusion models provide a powerful way to incorporate complex prior information for solving inverse problems. However, existing methods struggle to correctly incorporate guidance from conflicting signals in the prior and measurement, and often failed to maximizing the consistency to the measurement, especially in the challenging setting of non-Gaussian or unknown noise. To address these issues, we propose Measurement-Aligned Sampling (MAS), a novel framework for linear inverse problem solving that flexibly balances prior and measurement information. MAS unifies and extends existing approaches such as DDNM, TMPD, while generalizing to handle both known Gaussian noise and unknown or non-Gaussian noise types. Extensive experiments demonstrate that MAS consistently outperforms state-of-the-art methods across a variety of tasks, while maintaining relatively low computational cost.
@@ 1 Introduction
Here $\epsilon\in\mathbb{R}^{m}$} represents measurement noise, $x_{0}\in\mathbb{R}^{d}$ is drawn from data distribution $\pi_{0}(x_{0})$, $\mathcal{H}:\mathbb{R}^{d}\mapsto\mathbb{R}^{m}$ is the measurement function, and $y\in\mathbb{R}^{m}$ denotes the degraded measurement or observed image. A useful motivating example is a high-resolution image $x_{0}$, with a noisy degraded image $y$ and a known corruption process.
[…]
Moreover, the above approaches lack the ability to effectively handle unknown or non-Gaussian noise. In practical settings, noise frequently deviates from Gaussian assumptions—exhibiting characteristics like salt-and-pepper, periodic, or Poisson distributions—or is completely unknown. Additionally, the forward measurement operator may be uncertain or inaccurately specified. Effectively addressing inverse problems under these more general and realistic conditions remains an open and challenging research area.
@@
  • We propose Measurement-Aligned Sampling (MAS), a novel framework for solving linear inverse problems. MAS provides both probabilistic and optimization perspectives and generalizes approaches such as DDNM and TMPD for linear inverse problems. Furthermore, our proposed ‘overshooting’ technique achieves superior restoration quality compared to DDNM across various inverse problem scenarios.
  • We develop new techniques that maximize consistency with the measurement, enabling robust handling of both Gaussian noise and unknown noise sources. Moreover, our novel parameterization scheme allows us to effectively handle noisy inverse problems with unknown or non-Gaussian noise structures and even non-differentiable measurements, such as JPEG restoration, without requiring explicit knowledge of the forward operator or noise level. The comparison of method applicability across different inverse problems is shown in Table˜1.
  • Our experiments show that MAS enables robust and efficient image restoration, consistently outperforming baselines across Gaussian, non-Gaussian, and non-differentiable degradations (see Fig.˜1 and experiments in Sec.˜5).
@@
[…]
@@ 2 Background
Given training dataset $\mathcal{D}=\{x_{0}^{i}\}_{i=1}^{N}$ from target distribution $\pi_{0}(x_{0})$, $x_{0}^{i}\in\mathbb{R}^{d}$, the goal of generative modeling is to draw new samples from $\pi_{0}$. In the context of conditional generation, suppose that we have data samples from a joint distribution $(x_{0}^{i},y)\sim\pi(x_{0},y)$, where $x_{0}$, $y$ are dependent, and $y$ could be class labels or text information, for example.
For conditional generative modeling, we seek to draw new samples from $\pi(x_{0}\mid y)$ for a given condition $y$. Conditioned flows (zheng2023guided, ) build a marginal probability path $p_{t\mid y}$ using a mixture of interpolating densities: $p_{t\mid y}(x_{t}\mid y)=\int p_{t}(x_{t}\mid x_{0})\pi(x_{0}\mid y)dx_{T}$, where $p_{t}(\cdot\mid x_{0})$ is a probability path interpolating between noise and a single data point $x_{T}$. In general, the conditional kernel $p_{t}(x_{t}\mid x_{0})$ is given by a Gaussian distribution: $p_{t}(x_{t}\mid x_{0})=\mathcal{N}(x_{t};\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$, where $\mathcal{N}$ is the Gaussian kernel, $\alpha_{t},\sigma_{t}$ are differentiable functions. Then we can sample from the conditional distribution $p_{0\mid y}(x_{0}\mid y)$ by simulating a stochastic process $p_{t\mid y}(x_{t}\mid y)$ from time $t=T$ to $t=0$. Although different sampling methods can be chosen, generally, the iteration follows the form:
$x_{t-\Delta t}\sim\mathcal{N}(a_{t}m_{0\mid t,y}+b_{t}x_{t},c_{t}^{2}\mathbb{I}).$ (2)
where $m_{0\mid t,y}=\mathbb{E}[x_{0}\mid x_{t},y]$ is the idea conditional denoiser, $a_{t}$, $b_{t}$ and $c_{t}$ are parameters that depends on samplers. For instance, $x_{t-\Delta t}\sim\mathcal{N}(\alpha_{t-\Delta t}m_{0\mid t,y},\sigma_{t-\Delta t}\mathbb{I})$ is a valid DDIM sampler. In the implementation of conditional diffusion models, a denoiser is trained to approximate $m_{0\mid t,y}$. However, when only an unconditional denoiser $m_{0\mid t}=\mathbb{E}[x_{0}\mid x_{t}]$ is available, training-free conditional inference methods are employed.
Diffusion Posterior Sampling (DPS) and its variants. Given unconditional denoiser $\mathbb{E}[x_{0}\mid x_{t}]$, training-free conditional inference methods enable the approximation of the ideal conditional denoiser $\mathbb{E}[x_{0}\mid x_{t},y]$ (pokle2023training, ):
$\mathbb{E}[x_{0}\mid x_{t},y]=\mathbb{E}[x_{0}\mid x_{t}]+\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}\log p(y\mid x_{t}).$ (3)
Since $\nabla_{x_{t}}\log p(y\mid x_{t})$ is generally intractable, various approaches have been developed to approximate it.
For linear inverse problems, where the forward model is given by: $y=Hx_{0}+\epsilon$, $\epsilon\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I})$. Tweedie Moment Projected Diffusion (TMPD) (boys2023tweedie, ) assume $p(x_{0}\mid x_{t})$ as a Gaussian: $p(x_{0}\mid x_{t})\approx\mathcal{N}(m_{0\mid t},C_{0\mid t})$, where $m_{0\mid t}(x_{t}):=\mathbb{E}[x_{0}\mid x_{t}]$ is the ideal unconditional denoiser, $C_{0\mid t}(x_{t}):=\mathbb{E}[(x_{0}-m_{0\mid t})(x_{0}-m_{0\mid t})^{T}\mid x_{t}]$ is the covariance of $x_{0}\mid x_{t}$. Then the posteroir mean $\mathbb{E}[x_{0}\mid x_{t},y]$ admits an explicit closed-form solution:
$\mathbb{E}[x_{0}\mid x_{t},y]=m_{0\mid t}+C_{0\mid t}H^{T}(HC_{0\mid t}H^{T}+\sigma_{y}^{2}\mathbb{I})^{-1}(y-Hm_{0\mid t})$ (4)
The covariance $C_{0\mid t}$ could be calculated via gradient go through the denoiser: $C_{0\mid t}=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{m_{0\mid t}}(x_{t})$.
Optimization based methods. Unlike DPS guarantees that sampling is strictly from the conditional distribution, $p(x_{0}\mid y)$, optimization-based approaches (zhu2023denoising, ; li2024decoupled, ; wang2024dmplug, ) place more emphasis on the alignment with the measurement and the prior, which takes the following iteration:
$x_{0}^{*}$ $=\arg\min_{x_{0}}\;\bigl\|x_{0}-m_{0\mid t}\bigr\|^{2}+\lambda_{t}\,\bigl\|y-\mathcal{H}(x_{0})\bigr\|^{2},$ (5a)
$x_{t-\Delta t}$ $\sim\mathcal{N}\!\left(a_{t}x_{0}^{*}+b_{t}x_{t},\,c_{t}^{2}\right).$ (5b)
where $\lambda_{t}$ is a manually designed hyperparameter and $\mathcal{H}(\cdot)$ is the nonlinear forward operator. The iteration of optimization based methods could be seen as replacing $m_{0\mid t,y}$ in Eq.˜2 to $x_{0}^{*}$ in Eq.˜5a.
3 Methodology
For optimization based methods, the data-consistency loss with respect to the measurement $y$ is treated uniformly across all directions of the measurement space. However, for inverse problems it is often advantageous to introduce a weighting matrix that reflects the geometry of the forward operator (tarantola2005inverse, ). To this end, we propose Measurement-Aligned Sampling (MAS), which incorporates such a weighting into the optimization. As we demonstrate in Sec.˜5, this alignment leads to significant improvements in reconstruction quality.
3.1 Measurement Aligned Sampling
In this work, we generalize the objective in Eq.˜5a as
$x_{0}^{*}$ $=\arg\min_{x_{0}}\|x_{0}-m_{0\mid t}\|^{2}+\|y-Hx_{0}\|_{W^{-1}}^{2}.$ (6)
where $W\coloneqq\eta_{1}HH^{\mathsf{T}}+\eta_{2}\mathbb{I}$ (with $\eta_{1}\geq 0,\eta_{2}\geq 0$) is the weighted matrix and serves as a metric that balances measurement fidelity and prior regularization, where $\|z\|^{2}_{A}=z^{T}Az$.
When $\eta_{1}=0$ and $\eta_{2}>0$, corresponding to the classical Tikhonov (ridge) regularization, where $\eta_{2}$ controls the trade-off between fitting the measurements $y$ and staying close to the prior $m_{0\mid t}$. When $\eta_{1}>0$ and $\eta_{2}=0$, the data term becomes weighted by $(HH^{\top})^{-1}$, a Mahalanobis-type distance that emphasizes alignment along directions where $H$ is weak (small singular values), thereby regularizing ill-posed components of the inverse problem. The balance between $\eta_{1}$ and $\eta_{2}$ plays a crucial role in reconstruction quality, as we show in our experiments.
Finally, Eq.˜6 admits a unique closed-form solution obtained by setting the gradient to zero:
$x_{0}^{*}$ $=Y^{-1}[m_{0\mid t}+H^{\mathsf{T}}W^{-1}y]$ (7)
$\text{where}\quad W$ $\coloneqq\eta_{1}HH^{\mathsf{T}}+\eta_{2}\mathbb{I},\qquad Y\coloneqq\mathbb{I}+H^{\mathsf{T}}W^{-1}H,$
In practice, computing the inverse $W^{-1}$ and $Y^{-1}$ in Eq.˜7 naively can be computationally expensive. Instead, one can employ singular value decomposition (SVD) for more efficient computation; see Sec.˜B.2 for details.
Remark 1 (Connection with DDNM (wang2022zero, )).
As $\eta_{2}=0$ and $\eta_{1}\rightarrow 0$, $x_{0}^{*}\rightarrow\tilde{x}_{0}^{\text{DDNM}}:=m_{0\mid t}+H^{\dagger}(y-Hm_{0\mid t})$. Thus, in this limiting case, MAS recovers DDNM.
Remark 2 (Connection with optimization methods).
For the case where $\eta_{1}=0,\eta_{2}>0$, Eq.˜6 reproduces optimization approaches, such as Resample (song2023solving, ), DiffPIR (zhu2023denoising, ), DCDP (li2024decoupled, ), DMPlug (wang2024dmplug, ).
3.2 Probabilistic interpretation
We can interpret $x_{0}^{*}$ in Eq.˜7 as $\mathbb{E}[x_{\epsilon}\mid x_{t},y]$, where $x_{\epsilon}\approx x_{0}$ with perturbation variance $\sigma_{\epsilon}^{2}$ chosen to be sufficiently small so that $p(x_{0}\mid x_{\epsilon})\approx\mathcal{N}(x_{\epsilon},\sigma_{\epsilon}^{2}\mathbb{I})$. This formulation introduces an additional hyperprior, which—as demonstrated in our experiments—proves beneficial in addressing inverse problems.
Algorithm 1 Measurement-Aligned Sampling (MAS) for inverse problems.
1:Input: measurement $y$, forward operator $H(\cdot)$, pretrained DM $\epsilon_{\theta}(\cdot)$, number of diffusion step NN, diffusion schedule $\alpha_{t}$ and $\sigma_{t}$, objective parameters $\eta_{1}$, $\eta_{2}$.
2:Initialization: $x_{N}\sim\mathcal{N}(0,\mathbb{I})$
3:for $n=N$ to 11 do
4:  $\hat{x}_{0}\leftarrow[{x_{n}-\sigma_{n}\epsilon_{\theta}(x_{n},n)}]/{\alpha_{n}}$ $\triangleright$ Obtain predicted data $\mathbb{E}[x_{0}\mid x_{n}]$
5:  $x_{0}^{\prime}=Y^{-1}[\hat{x}_{0}+H^{\mathsf{T}}W^{-1}y]$ $\triangleright$ Calculating posterior mean $\mathbb{E}[x_{\epsilon}\mid x_{n},y]$
6:  $x_{n-1}\sim\mathcal{N}(\alpha_{n-1}x_{0}^{\prime},\sigma_{n-1}\mathbb{I})$ $\triangleright$ Forward diffusion step
7:end for
8:Output $x_{0}$
Given the measurement model $p(y\mid x_{0})=\mathcal{N}(Hx_{0},\sigma_{y}^{2}\mathbb{I})$ and conditional $p(x_{0}\mid x_{\epsilon})=\mathcal{N}(x_{\epsilon},\sigma_{\epsilon}^{2}\mathbb{I})$, the induced distribution over the measurement conditioned on $x_{\epsilon}$ takes the explicit form:
$p(y\mid x_{\epsilon})=\mathcal{N}\!\left(Hx_{\epsilon},\,\sigma_{y}^{2}\mathbb{I}+\sigma_{\epsilon}^{2}HH^{\mathsf{T}}\right).$ (8)
Notably, the likelihood $p(y\mid x_{\epsilon})$ shares the same mean as $p(y\mid x_{0})$, but with a generalized variance inflated by a term depending on $H$. Since both $p(y\mid x_{\epsilon})$ and $p(y\mid x_{0})$ are Gaussian, the posterior distribution admits a closed-form expression. In particular, the posterior mean $\mathbb{E}[x_{0}\mid x_{t},y]$ can be computed via Bayesian linear regression, as stated in Prop.˜3.1.
Proposition 3.1 (Bayesian Linear Regression).
Suppose $p(y\mid x_{\epsilon})=\mathcal{N}\!\left(Hx_{\epsilon},\,R\right)$, $R\coloneqq\sigma_{y}^{2}\mathbb{I}+\sigma_{\epsilon}^{2}HH^{\mathsf{T}}$ and $p(x_{\epsilon}\mid x_{t})\approx\mathcal{N}\!\left(m_{0\mid t},\,C_{0\mid t}\right)$. Then the posterior is Gaussian, with mean given by
$\mathbb{E}[x_{\epsilon}\mid x_{t},y]=\Big(C_{0\mid t}^{-1}+H^{\mathsf{T}}R^{-1}H\Big)^{-1}\Big(C_{0\mid t}^{-1}m_{0\mid t}+H^{\mathsf{T}}R^{-1}y\Big).$ (9)
As we set $C_{0\mid t}=r_{t}^{2}\mathbb{I}$, $\eta_{1}\coloneqq{\sigma_{\epsilon}^{2}}/{r_{t}^{2}}$ and $\eta_{2}\coloneqq\sigma_{y}^{2}/r_{t}^{2}$, $\mathbb{E}[x_{\epsilon}\mid x_{t},y]$ in Eq.˜9 is equivalent to $x_{0}^{*}$ in Eq.˜7, which provides a probabilistic perspective for MAS.
Remark 3 (Connection to TMPD (boys2023tweedie, )).
Setting $\sigma_{\epsilon}=0$ reduces the posterior mean in Eq.˜9 to that of TMPD in Eq.˜4.
Remark 4 (‘Overshooting’ trick).
Theoretically, $\eta_{1}\geq 0$ since $\eta_{1}\coloneqq\sigma_{\epsilon}^{2}/r_{t}^{2}$, however, the posterior mean in Eq.˜7 allows negative $\eta_{1}$. As illustrated in Fig.˜2, negative $\eta_{1}$ produces an overshooting effect, drawing $x_{0}^{*}$ even further toward alignment with the measurement $y$ than prescribed by DDNM. Interestingly, in our experiments this overshooting effect leads to improved reconstruction quality. A more detailed discussion is provided in Sec.˜B.2.
Figure 2: 2D illustration of the influence of parameters $\eta_{1}$ and $\eta_{2}$. Dots represent $x_{0}^{*}$, calculated via Eq.˜7. (a) Parameter $\eta_{1}$ controls the trade-off between $m_{0\mid t}$ and $\tilde{x}_{0}^{\text{DDNM}}$: as $\eta_{1}\rightarrow\infty$, the posterior mean $x_{0}^{*}$ approaches $m_{0\mid t}$; as $\eta_{1}\rightarrow 0$, it converges to $\tilde{x}_{0}^{\text{DDNM}}$. (b) Adjusting $\eta_{2}$ alters the posterior trajectory differently from varying $\eta_{1}$.
4 Maximizing the consistency for noisy inverse problems
4.1 Why previous methods failed to maximize the consistency?
DDNM highlighted that calculating the posterior sampling $\tilde{x}_{0}^{\text{DDNM}}=m_{0\mid t}+H^{\dagger}(y-Hm_{0\mid t})$ can inadvertently introduce additional noise into $x_{t}$, if $y$ is noisy. For instance, consider a simple forward model: $y=x_{0}+\epsilon_{y}$, where both $H$ and $H^{\dagger}$ are identity matrix, i.e., $H=H^{\dagger}=\mathbb{I}$, then $\tilde{x}_{0}^{\text{DDNM}}=y=x_{0}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\epsilon_{y}}$. Here $\epsilon_{y}$ is the additional noise introduced to $\tilde{x}_{0}^{\text{DDNM}}$, and will be further introduced into $x_{t-\Delta t}$. We argue that this issue is not unique to DDNM, but may also arise in TMPD (boys2023tweedie, ), DAPS (zhang2024improving, ), as well as in optimization-based methods (zhu2023denoising, ; li2024decoupled, ).
For MAS and under this same example, $y=x_{0}+\epsilon_{y}$, calculating $x_{0}^{*}$ (Eq.˜7) yields
$x_{0}^{*}=m_{0\mid t}+\frac{y-m_{0\mid t}}{\eta_{1}+\eta_{2}+1}=m_{0\mid t}+\frac{x_{0}-m_{0\mid t}}{\eta_{1}+\eta_{2}+1}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\epsilon_{y}}{\eta_{1}+\eta_{2}+1}},$ (10)
where $\epsilon_{y}/(\eta_{1}+\eta_{2}+1)$ is the additional noise introduced to $x_{0}^{*}$. A delicate balance arises from the fact that increasing either $\eta_{1}$ or $\eta_{2}$ will not only reduce the influence of the (unknown) noise term $\epsilon_{y}$, but also reduce the consistency with the measurement $y$ in general. To address this issue, we propose two approaches for addressing known Gaussian noise (Sec.˜4.2) and unknown noise Sec.˜4.3.
4.2 Addressing Gaussian noise with known variance
To handle Gaussian noise with known variance and $H=\mathbb{I}$, we modify Eq.˜10 and Eq.˜2 as:
$x_{0}^{*}=m_{0\mid t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda_{t}}\frac{y-m_{0\mid t}}{\eta_{1}+\eta_{2}+1},\quad x_{t-\Delta t}\sim\mathcal{N}(a_{t}\tilde{x}_{0}+b_{t}x_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t}}\mathbb{I}).$ (11)
Here $\lambda_{t}$ and $\gamma_{t}$ are two parameters that can control the total noise introduced to $x_{t-\Delta t}$. In our work, we adopt similar two principles as DDNM++ (wang2022zero, ) for handling Gaussian noise: (i) the total noise introduced in $x_{t-\Delta t}$ should be $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$ to conform to the correct distribution of $x_{t-\Delta t}$ in Eq.˜2; (ii) $\lambda_{t}$ should be as close to $1$ as possible to maximize the preservation of $x_{0}^{*}$. As $\epsilon_{y}\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I})$, principle (i) and principle (ii) are equivalent to: